40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4893     Entry time: Tue Jun 28 02:11:47 2011
Author: Jenne 
Type: Update 
Category: Locking 
Subject: Latest MICH noise budget 

I have measured / calculated the latest MICH noise budget.  It doesn't really look all that stellar.

MICH_noise_budget_as_of_28June2011.png

As you can see, we are nowhere near being shot noise limited, since there's a huge discrepancy between all of the measured spectra and the teal Shot Noise line. 

One possible suspect is that the analog whitening filters weren't on when I took my measurements.  I didn't actually check to ensure that they were on, so they might not have been.  Right now we're limited by electronics and other boring noises, so I need to make sure we're limited by the noise of the diode itself (we don't have enough light in the IFO to actually be shot noise limited since that takes 2.5mA for AS55 and I only have 1.1mA, but we should be ~within a factor of 2ish).

ELOG V3.1.3-